A standard deck of cards contains 52 cards. One card is selected from the deck.
​(a)
Compute the probability of randomly selecting a spade or diamond.
​(b)
Compute the probability of randomly selecting a spade or diamond or heart.
​(c)
Compute the probability of randomly selecting a five or club.

Answer by t0hierry(194)

There are 26 spade or diamond cards (out of 52), so the probability you draw one of these is 26/52 = 1/2
b) same answer 1/2
c) There are 4 fives and 13 clubs one of which is a five, so the number of distinct possibilities is 16. Divide by 52 and the probability is 16/52 = 4/13